# Misspecification uncertainties in near-deterministic regression

### v2

@@ Appendix D POPS-constrained variational minimization of ensemble ansatz
where we use integration by parts. Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models. An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix E.
@@
In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in section II.3. In this regime, for small $\sigma$ the variation of the upper bound (7) has a limiting form of
@@
in agreement with the result given in equation (21).
Appendix E POPS-constrained variational ansatz for PAC-Bayesian linear regression
The POPS-constrained ensemble ansatz can also be used to incorporate both epistemic and misspecification parameter uncertainties in a variational scheme robust at small $N/P$, which we detail here for completeness. We assume a Gaussian posterior $\pi_{V}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V},{\bm{\Sigma}}_{V})$, such that the ensemble ansatz is characterized by the ensemble mean ${\bm{\mu}}_{E}=\langle{\bm{\Theta}}^{*}_{\mathbf{X}}\rangle$ and covariance ${\bm{\Sigma}}_{E}$. A first approximation simply combines the covariance from minimization of the PAC-Bayes loss bound (7), ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\left[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}+N\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle_{\mathcal{X}}\right]^{-1}$, with the ensemble covariance ${\bm{\Sigma}}_{E}$. This yield a posterior covariance ${\bm{\Sigma}}^{*}_{\mathcal{L}}+{\bm{\Sigma}}_{E}$, which remains robust at finite $N/P$ and limits to ${\bm{\Sigma}}_{E}$ as $N/P\to\infty$.
However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models (Tipping, 2001). Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (23) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
${\bm{\mu}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}}+\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}{\bm{\Theta}}_{V},\quad{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]^{\top}\right\rangle_{\mathcal{X}},$ (34)
where ${\bf v}_{\mathbf{X}},{\bf W}_{\mathbf{X}}$ are derived from (23) such that ${\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\mu}}_{E}$ and ${\bm{\Sigma}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\Sigma}}_{E}$, reading
${\bf v}_{\mathbf{X}}\equiv{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad{\bf W}_{\mathbf{X}}\equiv\mathbb{I}-{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathbf{F}}^{\top}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad\delta{\bf v}_{\mathbf{X}}\equiv{\bf v}_{\mathbf{X}}-\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}},\quad\delta{\bf W}_{\mathbf{X}}\equiv{\bf W}_{\mathbf{X}}-\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}.$
This gives a variational distribution which can be used in the PAC-Bayes loss bound ^$\hat{{\mathcal{L}}}_{N}[\pi]$ (7)
$\pi_{V}({\bm{\Theta}}|{\bm{\Theta}}_{V})\equiv\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V}({\bm{\Theta}}_{V}),{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})),\quad{\bm{\Theta}}_{V}^{*}\equiv\arg\min_{{\bm{\Theta}}_{V}\in{\mathcal{P}}}\hat{{\mathcal{L}}}_{N}[\pi_{V}(.|{\bm{\Theta}}_{V})].$ (35)
Using the empirical loss $\hat{L}_{N}({\bm{\Theta}})$, quadratic in ${\bm{\Theta}}$ for linear models, the variational minimum satisfies
$\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}[N{\bm{\nabla}}_{\bm{\Theta}}\hat{L}_{N}({\bm{\Theta}}^{*}_{V})+{\bm{\Sigma}}_{0}^{-1}{\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{V})]+\left[N{\bf A}^{-1}+{\bm{\Sigma}}_{0}^{-1}+{\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}^{*}_{V})\right]:{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})={\bf 0},$ (36)
where :::: denotes a contraction over the first two indicies of a third-order tensor ${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})$ linear in ${\bm{\Theta}}$:
${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})=\left\langle\delta{\bf W}_{\mathbf{X}}\delta{\bf v}^{\top}_{\mathbf{X}}+\delta{\bf v}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\right\rangle_{\mathcal{X}}.$ (37)
As for Bayesian ridge regression, the only non-linearity in (36) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g. (Tipping, 2001),(Wipf and Nagarajan, 2007). A full study of the low data regime is left for future work.
Appendix F Implementation of the ensemble ansatz in Python
The following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}_{E}({\bf X})=1$:
@@ Appendix F Implementation of the ensemble ansatz in Python
4    # Inverse of A
5    A_i = np.dot(X.T,X) +
@@
15    return T
An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized.
Appendix G Additional plots of test error prediction
The plots below are identical in presentation to the top row of figure 3, but for the quadratic and random linear simulation engines described in the main text.
Figure 5: Test error prediction for medium dimensional problems. From left: Observed test errors for a misspecified linear surrogate model (P=20) on a quadratic simulation engine, along with predicted test errors from the POPS-ensemble. Parity plots for the minimum loss model, along with envelope bounds from the POPS-ensemble.

### v3

@@ D.1 Variational optimization
where we use integration by parts. Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models. An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix LABEL:app:variational-gaussian.
@@
In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in the main text. In this regime, for small $\sigma$ the variation of the upper bound (21) has a limiting form of
@@
in agreement with the result given in the main text.
Appendix E Implementation of the ansatz in Python
The hypercube projector ${\bf V}\in\mathbb{R}^{R\times P}$ and bounds ${\bf B}\in\mathbb{R}^{R\times 2}$ can be produced with the following python code, where A.T is the transpose of A:
@@ Appendix E Implementation of the ansatz in Python
4 # Inverse of A with prior matrix
5 A_i = np.dot(X.T,X) + prior / y.size
@@
15 # Find hypercube projector
16 nu,V = np.linalg.eigh(T@T.T)
17 nu_sel = nu<nu.max()*thresh
18 V = V[nu_sel,:] # shape = (R,P)
19 # bounds
20 B = np.array([(T@V.T).max(0),(T@V.T).min(0)]).T
21 return V,B
To sample the ensemble we use uniform random numbers
1import numpy as np
2def resample(V,B,n=20,seed=123):
3 np.random.seed(seed)
4 u = np.random.uniform(\
5 min=B[:,0],
6 max=B[:,1],
7 size=(n,V.shape[0]))
8 return u@V
Appendix F Additional plots of test error prediction
F.1 Low dimensional models
We applied the same procedure as for the polynomical example to a model system with $P=6$ independent features to simulate more realistic datasets. As before, the numerical minimization routine could only treat a small range of $\sigma$ values. Figure 4 shows the resulting parameter distributions. At large $\sigma=2$ the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$. In addition, a minimizer initialized with the POPS-ensemble did not change appreciably. These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
Figure 4: Regression of a linear model onto a cubic ”simulation engine” with $P=6$ independent features, trained on $N=500$ points. The panels show the cumulative distribution of each parameter across the $N$-member ensemble. Black, gray : ensemble and hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$. Blue, orange and green : numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=2,1,1/2$. As in Figure 1 in the main text, lower $\sigma$ gave numerical instabilities.
F.2 Medium dimensional models
The plots below are identical in presentation to the top row of figures in the main text, but for the quadratic and random linear simulation engines described in the main text.
Figure 5: Test error prediction for medium dimensional problems. Left: Observed test errors for a misspecified linear surrogate model (P=15) on a quadratic (top) and random linear (bottom) simulation engine, along with predicted test errors from the POPS-ensemble. Right: Parity plots for the minimum loss model, along with envelope bounds from the POPS-ensemble and hypercube.